# On the enumeration of connected sets in finite lattice graphs

### v1

@@ 4 Asymptotic analysis
    In fact, both bounds in (i) and (ii) provide a good estimate, as shown in Figure 5(a) and Table 3 that illustrate the asymptotic behavior of $c(P_{3}\times P_{n})$ and $c_{3}=(N_{L}(P_{3}\times P_{n}))^{1/3n}=\sqrt[3n]{6n+\sum\limits_{k=2}^{n}(\mathbf{1}^{T}A_{3}^{k-1}\mathbf{1})\cdot(n-k+1)}$, and Figure 5(b) and Table 4 show the asymptotic behavior of $c(K_{1,3}\times P_{n})$ and $c_{4}=(N_{L}(K_{1,3}\times P_{n}))^{1/4n}=\sqrt[4n]{11n+\sum\limits_{k=2}^{n}(\mathbf{1}^{T}A_{4}^{k-1}\mathbf{1})\cdot(n-k+1)}$, when $1\leq n\leq 800$.
@@ Acknowledgement
This work was supported by National Natural Science Foundation of China (No. 12171402).

### v2

@@ 4 Asymptotic analysis
    Moreover, $c_{3}=(N_{L}(P_{3}\times P_{n}))^{1/3n}$ and $c_{4}=(N_{L}(K_{1,3}\times P_{n}))^{1/4n}$ provide a good estimate of $c(P_{3}\times P_{n})$ and $c(K_{1,3}\times P_{n})$, respectively. See some data in Figure 6, Table 3 and Table 4 as examples that illustrate the asymptotic behavior of $c(P_{3}\times P_{n})$, $c(K_{1,3}\times P_{n})$, $c_{3}$ and $c_{4}$.
  4. (iv)
    Additionally, based on extensive Mathematica computations, we suspect that the limits of $c(P_{3}\times P_{n})$ and $c(K_{1,3}\times P_{n})$ exist as $n$ tends to infinity, with approximate values of 1.7196 and 1.7914, respectively.
@@ Acknowledgement
This work was supported by National Natural Science Foundation of China (No. 12571379 and 12501498).